# MUBen: Benchmarking the Uncertainty of Molecular Representation Models

### v3

@@ Abstract
Large molecular representation models pre-trained on massive unlabeled data have shown great success in predicting molecular properties. However, these models may tend to overfit the fine-tuning data, resulting in over-confident predictions on test data that fall outside of the training distribution. To address this issue, uncertainty quantification (UQ) methods can be used to improve the models’ calibration of predictions. Although many UQ approaches exist, not all of them lead to improved performance. While some studies have included UQ to improve molecular pre-trained models, the process of selecting suitable backbone and UQ methods for reliable molecular uncertainty estimation remains underexplored. To address this gap, we present MUBen, which evaluates different UQ methods for state-of-the-art backbone molecular representation models to investigate their capabilities. By fine-tuning various backbones using different molecular descriptors as inputs with UQ methods from different categories, we critically assess the influence of architectural decisions and training strategies on property prediction and uncertainty estimation. Our study offers insights for selecting UQ for backbone models, which can facilitate research on uncertainty-critical applications in fields such as materials science and drug discovery.
@@ 1 Introduction
Figure 1: MUBen overview with datasets, backbones, UQ methods and metrics enumerated.
We present MUBen, a benchmark designed to assess the performance of UQ methods applied to the molecular representation models for property prediction on various metrics, as illustrated in Figure 1. It encompasses UQ methods from different categories, including deterministic prediction, Bayesian neural networks, post-hoc calibration, and ensembles (§ 4.2), on top of a set of molecular representation (backbone) models, each relies on a molecular descriptor from a distinct perspective (§ 4.1). MUBen delivers intriguing results and insights that may guide the selection of backbone model and/or UQ methods in practice (§ 5). We structure our code, available at https://github.com/Yinghao-Li/MUBen, to be user-friendly, easily transferrable, and extendable, with the hope that this work will promote the future development of UQ methods, pre-trained models, or applications within the domains of materials science and drug discovery.
@@ 2.1 Pre-Training Molecular Representation Models
Existing pre-trained models for molecular representation learning can be distinguished into three categories based on the type of molecular descriptors.
  • The first category includes ChemBERTa (Chithrananda et al., 2020), SMILES-BERT (Wang et al., 2019), Molformer (Ross et al., 2022), GPT-MolBERTa (Balaji et al., 2023), etc., which leverage string-based input descriptors such as SMILES (Weininger, 1988), SELFIES (Krenn et al., 2020), Group SELFIES (Cheng et al., 2023), or SAFE (Noutahi et al., 2023). Such models are generally based on the Transformer encoder and pre-trained with masked language modeling (MLM) objectives (Devlin et al., 2019), which masks out a portion of tokens and trains the model to predict them using the context information.
  • Given that molecules can be represented as 2D topological graphs with atoms as vertices and chemical bonds as edges, the second category of pre-training strategies employs techniques such as context prediction (Hu et al., 2020; Rong et al., 2020), contrastive learning (Sun et al., 2020; You et al., 2020; Wang et al., 2022), and replaced component detection (Li et al., 2021; Kim et al., 2022) to learn molecular representation models with the awareness of the 2D atomic relationships.
@@
Beyond property prediction, pre-trained molecular representation models are leveraged for additional tasks, such as chemical design (Taylor et al., 2022; Ghugare et al., 2023; Li et al., 2023a). For instance, Ghugare et al. (2023) delve into the potential of merging text-based generative models with reinforcement learning to synthesize molecules with targeted properties. This trend further emphasize the value of MUBen.
@@ 2.2 Uncertainty Quantification
Uncertainty estimation plays a crucial role in stimulating deep neural networks to acknowledge what they don’t know, which has been applied to various domains such as object detection (Chen et al., 2021; He et al., 2019) and decision making (Amiri et al., 2020; Kong et al., 2022). UQ on neural networks generally takes two directions: Bayesian and non-Bayesian. Bayesian neural networks (BNNs) quantify predictive uncertainty by imposing probability distributions over model parameters instead of using point estimates. While BNNs provide a principled way of UQ, exact inference of parameter posteriors is often intractable. To circumvent this difficulty, existing works adopt Monte Carlo dropout (Gal & Ghahramani, 2016), variational inference (Blundell et al., 2015; Kingma et al., 2015) and Markov chain Monte Carlo (MCMC) (Welling & Teh, 2011; Maddox et al., 2019) for posterior approximation. Along the other direction, non-Bayesian methods estimate uncertainty by assembling multiple networks (Lakshminarayanan et al., 2017; Fort et al., 2019), adopting specialized output distribution (Sensoy et al., 2018; Amini et al., 2020; Soleimany et al., 2021) or training scheme (Lin et al., 2017; Mukhoti et al., 2020), or calibrating model outputs through additional network layers during validation and test (Platt et al., 1999; Guo et al., 2017; Song et al., 2019; Angelopoulos & Bates, 2021; Angelopoulos et al., 2021). Recently, there have been several UQ tools (Ghosh et al., 2021; Chung et al., 2021; Detommaso et al., 2023) and benchmarks from the general domain (Nado et al., 2021; Schmähling et al., 2023), biomedical engineering (Mehrtens et al., 2023; Greenman et al., 2022) and materials science (Hirschfeld et al., 2020; Hwang et al., 2020; Scalia et al., 2020; Varivoda et al., 2023; Wollschläger et al., 2023). MUBen also targets the last domain but covers a more diverse suite of tasks, stronger pre-trained molecular representation models, and a more comprehensive set of UQ methods. We hope MUBen can serve as a standard benchmark to faithfully evaluate the ability of UQ approaches.
@@ 3 Problem Setup
Let $\bm{x}$ denote a descriptor of a molecule, such as SMILES or a 2D topological graph. We define $y$ as the label with domain $y\in\{1,\ldots,K\}$ for $K$-class classification or $y\in$ for regression. $\bm{\theta}$ denotes all parameters of the molecular representation network and the task-specific layer. We assume that a training dataset $\mathcal{D}$ consists of $N$ i.i.d. samples $\mathcal{D}=\{(\bm{x}_{n},y_{n})\}_{n=1}^{N}$. Our goal is to fine-tune the parameters $\bm{\theta}$ on $\mathcal{D}$ and learn a calibrated predictive probability model $p_{\bm{\theta}}(y|\bm{x}_{n})$. The UQ metrics are briefly illustrated below with more details being provided in appendix D.
@@
  • Brier Score (BS, Brier, 1950) is the mean squared error between the predicted probability and the ground-truth labels, which serves as another proper scoring rule for classification tasks.
    $\mathrm{BS}=\frac{1}{N}\sum_{n=1}^{N}\sum_{y=1}^{K}(p_{\bm{\theta}}(y|\bm{x}_{n})-\bm{1}_{y=y_{n}})^{2},$
    where $\bm{1}$ is the indicator function, i.e., $\bm{1}_{a=b}=1$ if $a=b$; otherwhise $\bm{1}_{a=b}=0$.
  • Calibration Errors measure the correspondence between predicted probabilities and empirical accuracy. For classification, we use Expected Calibration Error (ECE, Naeini et al., 2015), which first divides the predicted probabilities $\{p_{\bm{\theta}}(y_{n}\mid\bm{x}_{n})\}_{n=1}^{N}$ into $S$ bins $B_{s}=\{n\in\{1,\dots,N\}\mid p_{\bm{\theta}}(y_{n}\mid\bm{x}_{n})\in(\rho_{s},\rho_{s+1}]\}$ with $s\in\{1,\dots,S\}$, each with size $N_{s}$, and then compute the $L_{1}$ loss between the accuracy and predicted probabilities within each bin:
    $\operatorname{ECE}=\sum_{s=1}^{S}\frac{N_{s}}{N}\left|\frac{1}{N_{s}}\sum_{n\in B_{s}}\bm{1}_{y_{n}=\hat{y}_{n}}-\frac{1}{N_{s}}\sum_{n\in B_{s}}p_{\bm{\theta}}(\hat{y}_{n}\mid\bm{x}_{n})\right|,$
@@
    Similarly, for regression, we use Regression Calibration Error (CE, Kuleshov et al., 2018), which calculates the true frequency of the predicted points lying in each confidence interval against the predicted fraction of points in that interval:
@@
    where $\frac{s}{S}$ is the expected quantile and $F_{\bm{\theta}}(y_{n};\bm{x}_{n})$ represents the predicted quantile value of $y_{n}$, e.g., Gaussian CDF $\Phi(y_{n};\hat{y}_{n},\hat{\sigma}_{n})$ parameterized by the predicted mean $\hat{y}_{n}$ and variance $\hat{\sigma}_{n}$ if we assume the labels are subject to the Gaussian distribution.
@@ 4.1 Backbone Models
For primary backbone models, we select various pre-trained Transformer encoder-based models, each is SOTA or close to SOTA for their respective input format: 1) ChemBERTa (Chithrananda et al., 2020; Ahmad et al., 2022), which accepts SMILES strings as input; 2) GROVER (Rong et al., 2020), pre-trained with 2D molecular graphs via a Transformer bolstered by a dynamic message-passing graph neural network; and 3) Uni-Mol (Zhou et al., 2023), which incorporates 3D molecular conformations into its input, and pre-trains specialized Transformer architectures to encode the invariant spatial positions of atoms and represent the edges between atom pairs. We also implement a 4) fully connected deep neural network (DNN) that uses $200$-dimensional RDKit features which are regarded as fixed outputs of a hand-crafted backbone feature generator. This simple model aims to highlight the performance difference between heuristic feature generators and the automatic counterparts that draw upon self-supervised learning processes for knowledge acquisition. In addition, we have two other backbones: 5) TorchMD-NET, as introduced by Thölke & Fabritiis (2022) and pre-trained according to Zaidi et al. (2023). This model is designed to handle 3D inputs and leverages an equivariant Transformer architecture aimed at quantum mechanical properties. 6) GIN from Xu et al. (2019), which processes 2D molecular graphs. GIN is theoretically among the most powerful graph neural networks (GNNs) available, offering an additional randomly initialized backbone baseline with less sophisticated input features. Compared with the primary one, these models are constrained in pre-training: TorchMD-NET is exclusively pre-trained on quantum mechanical data, whereas GIN does not involve any pre-processed features. Please refer to appendix B for more details.
@@ 4.2 Uncertainty Quantification Methods
In MUBen, we choose popular UQ methods from different categories. We provide a high-level sketch here of each method and leave the detailed description to appendix C.
@@
The ensemble method has long been used in machine learning to improve model performance (Dietterich, 2000) but was first adopted to estimate uncertainty in (Lakshminarayanan et al., 2017). It trains a deterministic network multiple times with different random seeds and combines their predictions at inference. The ensembles can explore training modes thoroughly in the loss landscape and thus are robust to noisy and out-of-distribution examples (Fort et al., 2019).
@@ 4.3 Datasets
In line with previous studies (Fang et al., 2022; Zhou et al., 2023), MUBen divides all datasets with scaffold splitting to minimize the impact of dataset randomness and, consequently, enhances the reliability of the evaluation. Moreover, scaffold splitting alienates the molecular features in each dataset split, inherently creating a challenging out-of-distribution (OOD) setup that better reflects the real-world scenario. For comparison, we also briefly report the results of random splitting. Please check appendix A for detailed dataset descriptions.
@@ 4.4 Training and Evaluation Protocols
We use Sigmoid output activation and binary cross-entropy (BCE) loss function on all classification models unless the UQ method specifies a different objective. For regression, we standardize the labels by mapping the training labels to a standard Gaussian distribution before model training and evaluation to mitigate the impact of label magnitude on multi-task datasets. During inference, the predicted values are converted back to their original distribution according to the label mean and variance from the training set before computing the metrics. For datasets with multiple prediction tasks such as QM9, a classification model features one output head for each task, while regression models have two—one for predicting mean $\hat{y}$ and the other for variance $\hat{\sigma}$, which is guaranteed to be positive using a SoftPlus activation (Lakshminarayanan et al., 2017). Regression models are trained with Gaussian NLL objective. We apply validation-based early stopping for parameter selection on all experiments unless the UQ method has other requirements.
@@
    We report the results from the best-performing UQ method—Deep Ensembles for TorchMD-NET and GIN. These backbones are not ranked together with the primary benchmark.
@@ 5 Results and Analysis
Tables 1 and 2 present the macro-averaged rankings of the primary backbone-UQ approaches across all datasets, as well as the scores on two representative datasets. Figure 2 provides further insights into primary backbone model performance by visualizing the Mean Reciprocal Ranks (MRRs) with averaged UQ reciprocal ranks. Compared to the ranks in the tables, MRRs accentuate top results, offering a complementary perspective on the backbone performance. Our analysis initiates with how the UQ methods contribute to the model prediction, followed by an exploration of the uncertainty quantification performance.
@@
Prediction Performance
Theoretically, inserting UQ methods into the training pipeline does not guarantee better prediction on i.i.d. datasets. However, since we ensure OOD test points with scaffold splitting, UQ methods may mitigate the distribution gap, yielding better test results. The columns ROC-AUC, RMSE, and MAE in Tables 1, 2, and Figure 2 illustrate the predictive performance of each method. Examining from the lens of UQ methods, none provides a consistent guarantee of performance improvement over direct prediction, except for Deep Ensembles, and MC Dropout for regression. The randomness in the initialization and training trajectory of Deep Ensembles explores a broader range of loss landscapes, which partially addresses the distribution shift issue, as observed in previous works (Garipov et al., 2018; Fort et al., 2019; Fang et al., 2023). MC Dropout samples multiple sets of model parameters from the Gaussian distributions centered at the fine-tuned deterministic net weights, which may flatten extreme regression abnormality triggered by OOD features. This phenomenon is less pronounced for classification due to the $(0,1)$ output domain. However, other BNNs do not exhibit the same advantage. SWAG, while similar to MC Dropout w.r.t. training, might intensify training data overfitting due to the additional steps taken to fit the parameter distribution. The stochastic sampling employed by BBP and SGLD complicates network training and may impact the prediction performance as a result.
@@
Figure 2: MRR of the backbone models, where each metric is macro-averaged from the reciprocal ranks of all corresponding UQ methods across all datasets. Uni-Mol stands out by consistently surpassing its counterparts in property prediction accuracy for both classification and regression tasks. Conversely, GROVER exhibits a more balanced performance in both prediction accuracy and UQ across various tasks.
Looking from the perspective of primary backbones, Uni-Mol secures the best prediction performance for both classification and regression. The superior molecular representation capability of Uni-Mol is attributed to the large network size, the various pre-training data and tasks, and the integration of results from different conformations of the same molecule. When contrasting DNN, ChemBERTa, and GROVER, it becomes apparent that the expressiveness of the molecular descriptors varies for different molecules/tasks. Interestingly, pre-trained models do not invariably surpass heuristic features when integrated with UQ methods. Selecting a model attuned to the specific task is advised over an indiscriminate reliance on “deep and complex” networks.
Uncertainty Quantification Performance
@@
Figure 3: The calibration plot of UQ methods. Here “true probability” refers to the empirical prediction accuracy, and “predicted probability” refers to the average predicted probability within each bin. These results reveal that Temperature Scaling typically enhances calibration effectiveness in the majority of scenarios, whereas Focal Loss tends to lead to over-calibration of the model, producing under-confident results.
Although limited in classification efficacy, both BBP and SGLD deliver commendable performance in predicting regression uncertainty, capturing $7$ out of $8$ top ranks for $4$ backbones on $2$ metrics, narrowly missing out to Deep Ensembles for the remaining one (Table 2). Yet, their inconsistent improvement of RMSE and MAE implies a greater influence on variance prediction than the mean. Figure 4 reveals SGLD’s tendency to “play safe” by predicting larger variances, while the deterministic method is prone to over-confident by ascribing small variances even to its inaccurate predictions. In addition, we do not observe a better correlation between SGLD’s error and variance. We assume that the noisy training trajectory prevents SGLD and BBP from sufficiently minimizing the gap between the predicted mean and true labels, thus encouraging them to maintain larger variances to compensate for the error. Please refer to § C.2 and § E.1 for more analysis on UQ performance.
@@
Figure 4: The absolute error between the model-predicted mean and true labels against the predicted standard deviation. We compare the performance of SGLD with the deterministic prediction on different backbones and datasets. The “$y=kx$” lines indicate whether the true labels lie within the $k$-std range of the predicted Gaussian. Also, a model is perfectly calibrated when its output points are arranged on an “$y=kx$” line for an arbitrary $k$. Notably, SGLD is observed to generate a larger variance for OOD samples, which tends to correspond more closely with the prediction errors.

### v4

@@ Abstract
Large molecular representation models pre-trained on massive unlabeled data have shown great success in predicting molecular properties. However, these models may tend to overfit the fine-tuning data, resulting in over-confident predictions on test data that fall outside of the training distribution. To address this issue, uncertainty quantification (UQ) methods can be used to improve the models’ calibration of predictions. Although many UQ approaches exist, not all of them lead to improved performance. While some studies have included UQ to improve molecular pre-trained models, the process of selecting suitable backbone and UQ methods for reliable molecular uncertainty estimation remains underexplored. To address this gap, we present MUBen, which evaluates different UQ methods for state-of-the-art backbone molecular representation models to investigate their capabilities. By fine-tuning various backbones using different molecular descriptors as inputs with UQ methods from different categories, we assess the influence of architectural decisions and training strategies on property prediction and uncertainty estimation. Our study offers insights for selecting UQ for backbone models, which can facilitate research on uncertainty-critical applications in fields such as materials science and drug discovery.
@@ 1 Introduction
Figure 1: MUBen’s pipeline with datasets, backbones, UQ methods and metrics enumerated.
We present MUBen, a benchmark designed to assess the performance of UQ methods applied to the molecular representation models for property prediction on various metrics, as illustrated in Figure 1. It encompasses UQ methods from different categories, including deterministic prediction, Bayesian neural networks, post-hoc calibration, and ensembles (§ 4.2), on top of a set of molecular representation (backbone) models, each relies on a molecular descriptor from a distinct perspective (§ 4.1). MUBen delivers intriguing results and insights that may guide the selection of backbone model and/or UQ methods in practice (§ 5). We structure our code, available at https://github.com/Yinghao-Li/MUBen, to be user-friendly and easily transferrable and extendable, with the hope that this work will promote the future development of UQ methods, pre-trained models, or applications within the domains of materials science and drug discovery.
@@ 2.1 Pre-Training Molecular Representation Models
Existing pre-trained models for molecular representation learning can be distinguished into three categories based on the type of molecular descriptors. The first category includes ChemBERTa (Chithrananda et al., 2020), SMILES-BERT (Wang et al., 2019), Molformer (Ross et al., 2022), GPT-MolBERTa (Balaji et al., 2023), etc., which leverage string-based input descriptors such as SMILES (Weininger, 1988), SELFIES (Krenn et al., 2020), Group SELFIES (Cheng et al., 2023), or SAFE (Noutahi et al., 2023). Such models are generally built upon the Transformer encoder and pre-trained with masked language modeling (MLM) objectives (Devlin et al., 2019), which masks out a portion of tokens and trains the model to predict them using the context information.
Given that molecules can be represented as 2D topological graphs with atoms as vertices and chemical bonds as edges, the second category of pre-training strategies employs techniques such as context prediction (Hu et al., 2020; Rong et al., 2020), contrastive learning (Sun et al., 2020; You et al., 2020; Wang et al., 2022), and replaced component detection (Li et al., 2021; Kim et al., 2022) to learn molecular representation models with the awareness of the 2D atomic relationships. Graph neural networks (GNNs, Scarselli et al., 2009) as well as Transformer encoders are commonly selected as the backbone architectures (Xia et al., 2023).
@@
Beyond property prediction, pre-trained molecular representation models are leveraged for additional tasks such as chemical design (Taylor et al., 2022; Ghugare et al., 2023; Li et al., 2023a). For instance, Ghugare et al. (2023) delve into the potential of merging text-based generative models with reinforcement learning to synthesize molecules with targeted properties. This trend further emphasize the value of MUBen.
@@ 2.2 Uncertainty Quantification
Uncertainty estimation plays a crucial role in stimulating deep neural networks to acknowledge what they don’t know, which has been applied to various domains such as object detection (He et al., 2019; Chen et al., 2021), decision making (Amiri et al., 2020; Kong et al., 2022), and text classification and information extraction (Chowdhury et al., 2011; Mukherjee & Awadallah, 2020; Yu et al., 2022; Li et al., 2022). UQ on neural networks generally takes two directions: Bayesian and non-Bayesian. Bayesian neural networks (BNNs) quantify predictive uncertainty by imposing probability distributions over model parameters instead of using point estimates. While BNNs provide a principled way of UQ, exact inference of parameter posteriors is often intractable. To circumvent this difficulty, existing works adopt Monte Carlo dropout (Gal & Ghahramani, 2016), variational inference (Blundell et al., 2015; Kingma et al., 2015) and Markov chain Monte Carlo (MCMC) (Welling & Teh, 2011; Maddox et al., 2019) for posterior approximation. Along the other direction, non-Bayesian methods estimate uncertainty by assembling multiple networks (Lakshminarayanan et al., 2017; Fort et al., 2019), adopting specialized output distribution (Sensoy et al., 2018; Amini et al., 2020; Soleimany et al., 2021) or training scheme (Lin et al., 2017; Mukhoti et al., 2020), or calibrating model outputs through additional network layers during validation and test (Platt et al., 1999; Guo et al., 2017; Song et al., 2019; Angelopoulos & Bates, 2021; Angelopoulos et al., 2021).
Recently, there have been several UQ tools (Ghosh et al., 2021; Chung et al., 2021; Detommaso et al., 2023) and benchmarks from the general domain (Nado et al., 2021; Schmähling et al., 2023), biomedical engineering (Mehrtens et al., 2023; Greenman et al., 2022) and materials science (Hirschfeld et al., 2020; Hwang et al., 2020; Scalia et al., 2020; Varivoda et al., 2023; Wollschläger et al., 2023). MUBen also targets the last domain but covers a more diverse suite of tasks, stronger pre-trained molecular representation models, and a more comprehensive set of UQ methods. We hope MUBen can serve as a standard benchmark to faithfully evaluate the ability of UQ approaches.
@@ 3 Problem Setup
Let $\bm{x}$ denote a descriptor of a molecule, such as SMILES or a 2D topological graph. Variable $y$ is the label with domain $y\in\{1,\ldots,K\}$ for $K$-class classification or $y\in$ for regression. $\bm{\theta}$ denotes all parameters of the molecular representation network and the task-specific layer. We assume that a training dataset $\mathbb{D}$ consists of $N$ i.i.d. samples $\mathbb{D}=\{(\bm{x}_{n},y_{n})\}_{n=1}^{N}$. Our goal is to fine-tune the parameters $\bm{\theta}$ on $\mathbb{D}$ in order to develop a predictive model that accurately estimates the probability distribution $p_{\bm{\theta}}(y|\bm{x}_{n})$. The UQ metrics are briefly illustrated below with more details being provided in appendix D.
@@
  • The Brier Score (BS, Brier, 1950) quantifies the mean squared error between predicted probabilities and actual labels, and is recognized as a proper scoring rule for classification tasks. The formula for the Brier Score is given by:
    $\mathrm{BS}=\frac{1}{N}\sum_{n=1}^{N}\sum_{y=1}^{K}\left(p_{\bm{\theta}}(y|\bm{x}_{n})-\mathbb{I}(y=y_{n})\right)^{2},$
    where $\mathbb{I}(\cdot)$ is the indicator function. Specifically, $\mathbb{I}(a=b)=1$ if $a=b$; otherwise, $\mathbb{I}(a=b)=0$.
  • Calibration Errors measure the correspondence between predicted probabilities and empirical accuracy. For classification, we use Expected Calibration Error (ECE, Naeini et al., 2015), which first divides the predicted probabilities $\{p_{\bm{\theta}}(y_{n}\mid\bm{x}_{n})\}_{n=1}^{N}$ into $S$ bins $\mathbb{B}_{s}=\{n\in\{1,\dots,N\}\mid p_{\bm{\theta}}(y_{n}\mid\bm{x}_{n})\in(\rho_{s},\rho_{s+1}]\}$ with $s\in\{1,\dots,S\}$, each with size $N_{s}$, and then compute the $L_{1}$ loss between the accuracy and predicted probabilities within each bin:
    $\operatorname{ECE}=\sum_{s=1}^{S}\frac{N_{s}}{N}\left|\frac{1}{N_{s}}\sum_{n\in\mathbb{B}_{s}}\mathbb{I}(y_{n}=\hat{y}_{n})-\frac{1}{N_{s}}\sum_{n\in\mathbb{B}_{s}}p_{\bm{\theta}}(\hat{y}_{n}\mid\bm{x}_{n})\right|,$
@@
    For regression tasks, the Regression Calibration Error (CE, Kuleshov et al., 2018) measures the accuracy of prediction intervals. It calculates the true frequency of the predicted points falling within each confidence interval against the predicted fraction of points in that interval:
@@
    where $\frac{s}{S}$ represents the expected quantile. $F_{\bm{\theta}}(y_{n};\bm{x}_{n})$ denotes the predicted quantile value for $y_{n}$, such as a Gaussian cumulative distribution function $\Phi(y_{n};\hat{y}_{n},\hat{\sigma}_{n})$, which is parameterized by the estimated mean $\hat{y}_{n}$ and variance $\hat{\sigma}_{n}$ assuming a Gaussian distribution for the labels.
@@ 4.1 Backbone Models
For primary backbone models, we select various pre-trained Transformer encoder-based models, each is SOTA or close to SOTA for their respective input format: 1) ChemBERTa (Chithrananda et al., 2020; Ahmad et al., 2022), which accepts SMILES strings as input; 2) GROVER (Rong et al., 2020), which pre-trains Transformer encoders augmented with the dynamic message-passing GNN on 2D molecular graphs ; and 3) Uni-Mol (Zhou et al., 2023), which incorporates 3D molecular conformations into its input, and pre-trains specialized Transformer architectures to encode the invariant spatial positions of atoms and represent the edges between atom pairs. We also implement a 4) fully connected deep neural network (DNN) that uses $200$-dimensional RDKit features which are regarded as fixed outputs of a hand-crafted backbone feature generator. This simple model aims to highlight the performance difference between heuristic feature generators and the automatic counterparts that draw upon self-supervised learning processes for knowledge acquisition. In addition, we have two other backbones: 5) TorchMD-NET, as introduced by Thölke & Fabritiis (2022) and pre-trained according to Zaidi et al. (2023). This model is designed to handle 3D inputs and leverages an equivariant Transformer architecture aimed at quantum mechanical properties. 6) GIN from Xu et al. (2019), which processes 2D molecular graphs. GIN is theoretically among the most powerful GNNs available, offering an additional randomly initialized backbone baseline with less sophisticated input features. Compared with the primary one, these models are constrained in pre-training: TorchMD-NET is exclusively pre-trained on quantum mechanical data, whereas GIN does not involve any pre-processed features. Please refer to appendix B for more detailed introduction on backbone architectures and hyperparameters.
@@ 4.2 Uncertainty Quantification Methods
In MUBen, we choose popular UQ methods from different categories. We provide a high-level sketch here of each method and leave the details to appendix C.
@@
The ensemble method has long been used in machine learning to improve model performance (Dietterich, 2000) but was first adopted to estimate uncertainty in (Lakshminarayanan et al., 2017). It trains a deterministic network multiple times with different random seeds and combines their predictions at inference. The ensembles can explore training modes thoroughly in the loss landscape and thus are robust to noisy and out-of-distribution (OOD) examples (Fort et al., 2019).
@@ 4.3 Datasets
In line with previous studies (Fang et al., 2022; Zhou et al., 2023), MUBen divides all datasets with scaffold splitting to minimize the impact of dataset randomness and, consequently, enhances the reliability of the evaluation. Moreover, scaffold splitting alienates the molecular features in each dataset split, inherently creating a challenging OOD setup that better reflects the real-world scenario. For comparison, we also briefly report the results of random splitting. Please check appendix A for detailed descriptions.
@@ 4.4 Training and Evaluation Protocols
We use Sigmoid output activation and binary cross-entropy (BCE) loss function on all classification models unless the UQ method specifies a different objective. For regression, we standardize the labels by mapping the training labels to a standard Gaussian distribution before model training and evaluation to mitigate the impact of label magnitude on multi-task datasets. During inference, the predicted values are converted back to their original distribution according to the label mean and variance from the training set before computing the metrics. The numbers of tasks are available in Table 4. A classification model features one output head for each prediction task, while regression models have two—one for predicting mean $\hat{y}$ and the other for variance $\hat{\sigma}$, which is guaranteed to be positive by applying a SoftPlus activation upon the output logits (Lakshminarayanan et al., 2017). Regression models are trained with Gaussian NLL objective. We universally apply validation-based early stopping unless the UQ method has other requirements.
@@
    We report the results from the best-performing UQ method—Deep Ensembles for TorchMD-NET and GIN. TorchMD-NET and GIN are not ranked together with the primary backbones while reporting the “Average Ranking”.
@@ 5 Results and Analysis
Tables 1 and 2 present the macro-averaged rankings of the primary backbone-UQ combinations across all datasets, as well as the metric scores on $4$ representative datasets, two for classification and two for regression. Figure 2 provides further insights into primary backbone model performance by visualizing the Mean Reciprocal Ranks (MRRs) with averaged across UQ methods. Compared to the ranks in the tables, MRRs accentuate top results, offering a complementary perspective on the backbone performance. Our analysis initiates with how the UQ methods contribute to the models’ property prediction, followed by an exploration of the uncertainty estimation performance.
@@
Property Prediction Performance
Theoretically, inserting UQ methods into the training pipeline does not guarantee better property prediction on datasets containing i.i.d. instances. However, since we ensure OOD test points with scaffold splitting, UQ methods may mitigate the distribution gap, yielding better test results. The columns ROC-AUC, RMSE, and MAE in Tables 1, 2, and Figure 2 illustrate the predictive performance of each method. Examining from the lens of UQ methods, none provides a consistent guarantee of performance improvement over the deterministic baseline, except for Deep Ensembles, and MC Dropout for regression. The randomness in the initialization and training trajectory of Deep Ensembles explores a broader range of loss landscapes, which partially addresses the distribution shift issue, as observed in previous works (Garipov et al., 2018; Fort et al., 2019; Fang et al., 2023). MC Dropout samples multiple sets of model parameters from the Gaussian distributions centered at the fine-tuned deterministic net weights, which may flatten extreme regression abnormality triggered by OOD features. This phenomenon is less pronounced for classification due to the $(0,1)$ output domain. However, other BNNs do not exhibit the same advantage. SWAG, while similar to MC Dropout w.r.t. training, might intensify training data overfitting due to the additional steps taken to fit the parameter distribution. While mitigating overfitting, the stochastic sampling methods used in BBP and SGLD can hinder network convergence, leading to insufficient exploration of the feature and loss spaces, and consequently, adversely affecting prediction performance. This impact is particularly acute in smaller models, which are inherently less expressive, as illustrated in Figure 10.
@@
Figure 2: MRR of the backbone models; each metric is macro-averaged from the reciprocal ranks of all corresponding UQ methods across all datasets. Uni-Mol stands out by consistently surpassing its counterparts in property prediction accuracy for both classification and regression tasks. Conversely, GROVER exhibits a more balanced performance in both prediction accuracy and UQ across various tasks.
Looking from the perspective of primary backbones, Uni-Mol secures the best prediction performance for both classification and regression. The superior molecular representation capability of Uni-Mol is attributed to the large network size, the various pre-training data and tasks, and the integration of results from different conformations of the same molecule (Zhou et al., 2023). When contrasting DNN, ChemBERTa, and GROVER, it becomes apparent that the expressiveness of the molecular descriptors varies for different molecules/tasks. Moreover, pre-trained models do not invariably surpass heuristic features when integrated with UQ methods, which aligns with the findings of Deng et al. (2023). Selecting a model attuned to the specific task is advised over an indiscriminate reliance on “deep and complex” networks.
@@
Figure 3: The calibration plot of UQ methods. Here, “true probability” refers to the empirical prediction accuracy, and “predicted probability” refers to the average predicted probability within each bin (§ 3). These results reveal that Temperature Scaling typically enhances calibration effectiveness in the majority of scenarios, whereas Focal Loss tends to lead to over-calibration of the model, producing under-confident results.
Uncertainty Quantification Performance
@@
Figure 4: The absolute error between the model-predicted mean and true labels against the predicted standard deviation. We compare the performance of SGLD with the deterministic prediction on different backbones and datasets. The “$y=kx$” lines indicate whether the true labels lie within the $k$-std range of the predicted Gaussian. Also, a model is perfectly calibrated when its output points are arranged on an “$y=kx$” line for an arbitrary $k$. Notably, SGLD is observed to generate a larger variance for OOD samples, which tends to correspond more closely with the prediction errors on average.
Although limited in classification efficacy, both BBP and SGLD deliver commendable performance in predicting regression uncertainty, capturing $7$ out of $8$ top ranks for $4$ backbones on two metrics, narrowly missing out to Deep Ensembles for the remaining one (Table 2). Yet, their inconsistent improvement of RMSE and MAE implies a greater influence on variance prediction than the mean. Figure 4 reveals SGLD’s tendency to “play safe” by predicting larger variances, while the deterministic method is prone to over-confident by ascribing small variances even to its inaccurate predictions. In addition, we do not observe a better correlation between SGLD’s error and variance. We assume that the noisy training trajectory prevents SGLD and BBP from sufficiently minimizing the gap between the predicted mean and true labels, thus encouraging them to maintain larger variances to compensate for the error. Please refer to § C.2 and § E.1 for more analysis on UQ performance.